OpenStudy (anonymous):

I Will Fan You And Medal You! Please Help! Compare and Contrast: Below are two expressions. Simplify each and then choose the statement that is true. Expression #1 (3d)3(d) Expression #2 (3d2)2 A.The exponents in Expression #1 are greater than the exponents of Expression #2. B.The exponents on Expression #2 are greater than the exponents of Expression #1. C.The exponents of Expression #1 are the same as the exponents of Expression #2. D.The relationship cannot be determined with the given information. I just need to know if the exponets add up to more in expression1or2

OpenStudy (anonymous):

@Mertsj Do you think you know this one

OpenStudy (anonymous):

@serena13579

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@Jhannybean @Jhannybean

OpenStudy (mertsj):

What is the first expression when you get it simplified?

OpenStudy (anonymous):

(3d)^3\[(3d)^3(d) and (3d^2)^2\]

OpenStudy (anonymous):

Is that what you ment @Mertsj

OpenStudy (mertsj):

What is the first one when simplified?

OpenStudy (anonymous):

I'm Not sure on How to simplify these :( Sorry

OpenStudy (mertsj):

\[(3d)^3(d)=(3d)(3d)(3d)(d)\]

OpenStudy (anonymous):

Oh that looks a lot more simpler

OpenStudy (mertsj):

\[(3d)(3d)(3d)(d)=(3)(3)(3)(d)(d)(d)(d)\]

OpenStudy (mertsj):

What is it?

OpenStudy (anonymous):

3^3 d^4

OpenStudy (mertsj):

3)(3)(3)=27 (d)(d)(d)(d)=d^4

OpenStudy (anonymous):

Oh , Sorry I'm A little slow :(

OpenStudy (mertsj):

So the first one is 27d^4

OpenStudy (mertsj):

Now you do the second one.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(3d^2)^2 (3d) (3d)^2

OpenStudy (anonymous):

(3) (3) (d) (d) ^2

OpenStudy (mertsj):

\[(3d^2)^2=(3d^2)(3d^2)=(3)(3)(d^2)(d^2)=(3)(3)(d)(d)(d)(d)\]

OpenStudy (mertsj):

Simplify

OpenStudy (mertsj):

What is (3)(3)?

OpenStudy (anonymous):

9d^4

OpenStudy (mertsj):

Now read your answer choices and see which one it is.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

The answer would be A

OpenStudy (anonymous):

No . NO my bad the answer is C @Mertsj

OpenStudy (mertsj):

Yes. They are the same.

OpenStudy (anonymous):

YAY! Thank You for not giving up on me it means a lot!

OpenStudy (mertsj):

yw